I am having trouble with inverse operations in having to make x the subject. I have three questions in my workbook.
First question - make x the subject.
v= qx^2/t+p
I tried doing it but I am not sure if I have my working right. My end answer was that it is
x= sqrt(v-pt/q)
Any help would be appreciated thank you. This is the one that I understood the most. The other two absolutely blow my mind. Thanks :)

Answer by Alan3354(69443)

First question - make x the subject.
v= qx^2/t+p
I tried doing it but I am not sure if I have my working right. My end answer was that it is x= sqrt(v-pt/q)
-------------
v= qx^2/t+p
Is the denominator t? or t+p?
Add parentheses.